OpenStudy (anonymous):

write y =3sin 8t−2.5cos 8t in the model form y= sqrt a^2 + b^2 sin (Bt+C)

sam (.sam.):

\[3\sin 8t−2.5\cos 8t=Rsin(\theta+\alpha)\]

OpenStudy (anonymous):

i dont understand